# Effect Identification and Unit Categorization in the Multi-Score Regression Discontinuity Design with Application to LED Manufacturing

### v1

@@ 3.3 Setup
We use following notation $\mathbb{R}^{K}_{>0}\coloneqq\left\{x\in\mathbb{R}^{K}\,\middle|\,\forall\,0\leq j\leq K\,:\,x_{j}>0\right\}$ for vectors with positive components and adapt it in the canonical way for the cross product of other ordered sets and relations (e.g. $\mathbb{R}^{K}_{<y}$ or $\mathbb{N}_{\leq k}$). Further, let $e_{1},\ldots,e_{K}\in\mathbb{R}^{K}$ denote the standard basis and $\left\langle\cdot,\,\cdot\right\rangle$ the standard scalar product in $\mathbb{R}^{K}$. For linear spaces $U$ and $V$ over $\mathbb{R}$ with $U\cap V=\{0\}$ we write $U\oplus V$ for their direct sum.
For each individual $i$, let $X_{i}=(X_{1,i},\,\ldots,\,X_{K,i})\in\mathbb{R}^{K}$ denote the score variables. Further, given $c\in\mathbb{R}^{K}$, let $I_{i}(c)\coloneqq(I_{1,i}(c_{1}),\,\ldots,\,I_{K,i}(c_{K}))$ denote the corresponding indicator variables with $I_{k,i}(c_{k})\coloneqq\operatorname{\mathbbm{1}}[X_{k,i}>c_{k}]$ and let the observed outcome for individual $i$ be $Y_{i}$. We regard the entries $I_{i}(c)$ as boolean variables and allow any composition of AND, OR and negation operations ($\land,\,\lor$$\overline{\square}$) over this set of atoms to form general boolean functions $g(I_{i}(c))$.
@@
A mapping $T:\mathbb{R}^{K}\to\{0,1\}$ is called a decision rule. We say that a decision rule $T$ is a cutoff rule (on $\mathbb{R}^{K}$) if there exists a boolean mapping $g$ such that $T_{i}=T(X_{i})=g(I_{i})$. That is $T$ explains the individual decisions $T_{i}$ using cutoffs according to $X_{i}$.
@@
holds for $\epsilon\in\mathbb{R}^{K},\lambda\in\mathbb{R}_{>0}$. Thus, without loss of generality, we assume that the cutoff of interest is $c=0$. We suppose that $Y_{i}$ depends in the following way on a cutoff rule $T$ and a general decision rule $D$:
@@
where $T$ is the treatment assignment and $D$ is the implemented treatment. Unless otherwise stated, we make no further assumptions on $D$ in the proofs, except that it is a decision-rule.
@@ 3.4 Unit Categorization
Given this setup, there are certain groups of individuals that are especially interesting, namely the nevertaker, the alwaystaker, the complier and the defier with respect to the pair (T,D)(T,\,D) or (G,D)(G,\,D), where $G$ is a subrule of $T$. We follow the intuition that a meaningful categorization of the behavior of a unit $i$ is only possible by means of counterfactual reasoning. That is, we need to investigate how imagined changes in the cutoff, or equivalently in the observed score values would have changed the responses $T_{i}$ and $D_{i}$.
@@
Figure 1: Cutoff rules $D=I_{1}\land I_{2}$ and $T=I_{1}$. The decision boundaries coincide with the coordinate axes. $T$ complies with $D$ given $X_{2}>0$. Otherwise, $D$ is zero independent of all variables that control the behavior of $T$, namely $X_{1}$.
Figure 1 illustrates this idea for $D\coloneqq I_{1}\land I_{2}$ and $T\coloneqq I_{1}$, and shows that only changes $c\in\mathbb{R}^{2}$ that impact $T$ are relevant to categorize the behavior of a unit $i$ regarding (T,D)(T,\,D). Additional variables that only affect $D$ might not be controllable or even known. The following definition formalizes the intuition behind relevant change directions of a cutoff rule.
@@
Let $T$ be a cutoff rule on $\mathbb{R}^{K}$. Denote with
@@
the support directions of $T$. The support of $T$ is defined as the linear hull over the support directions:
@@
This rather technical definition greatly simplifies the treatment of equivalent cutoff-rules e.g. $T\coloneqq I_{1}\land(I_{2}\lor\overline{I_{2}})$ and $G\coloneqq I_{1}$. Additionally, it allows comparing $T$ and $D$ on their common domain. While above notion of the support of a cutoff rule is motivated by the existence of a change along a single coordinate direction, it is almost trivial to see that an empty support also excludes changes along multiple directions. This justifies the definition.
@@
From now on, we require that $T$ is not degenerate in the sense that $\operatorname{supp}(T)\neq\{0\}$. For the introductory AND-rule example, $D\coloneqq I_{1}(c)\land I_{2}(c)$ and $T\coloneqq I_{1}(c)$ we have $\operatorname{supp}(T)=\mathbb{R}\times\{0\}$ and $\operatorname{supp}(D)=\mathbb{R}^{2}$. If a unit $i$ complies in the assigned treatment $T$ with the actual $D$, one would require that $T_{i}$ produces the same output as $D_{i}$, even under any hypothetical change of the cutoff $c_{1}$ as indicated in Figure 1. A more refined, but local definition could relax this requirement to “any reasonable changes”. For ease of presentation we stick to the global one. In other words, both rules should be synchronous on the support of $T$, which is the case if $X_{2,i}>c_{2}$. Thus, it is $X_{2,i}$ (or equivalently $c_{2}$) that controls the behavior of $i$. To make this distinction more apparent, it is useful to have a notation for entries $X\in\mathbb{R}^{K}$ that do not affect $T$:
@@
For each $X\in\mathbb{R}^{K}$, there exists a unique decomposition $X=X^{T}+X^{\perp T}$ with $X^{T}\in\operatorname{supp}(T)$ and $X^{\perp T}\in N^{T}$. The decomposition according to the orthogonal projection $P_{T}(X)\coloneqq\sum_{k\in S(T)}\left\langle X,\,e_{k}\right\rangle e_{k}$ onto $\operatorname{supp}(T)$ satisfies the above properties.
That is, according to Proposition 1 one has the following decomposition
$\mathcal{S}\coloneqq\mathbb{R}^{|S(T)|}\times\mathbb{R}^{K-|S(T)|}\simeq\operatorname{supp}(T)\,\oplus N^{T}=\mathbb{R}^{K}$
of the score space into two distinct parts. The former part captures the decision changes of $T$ and the latter consists of free variables not affecting $T$, but potentially controlling the unit category. This observation allows for a general definition of unit categories.
@@
Let $i$ be a unit and $X_{i}=X_{i}^{T}+X_{i}^{\perp T}$. Then $i$ is said to be
@@
for all $c\in\operatorname{supp}(T)$. Let the set of complier, nevertaker, alwaystaker and defier be denoted by $\operatorname{ComP}(T,D)$, $\operatorname{At}(T,D)$, $\operatorname{Nt}(T,D)$ and $\operatorname{DeF}(T,D)$.
Particularly, if $D$ is a cutoff-rule one obtains the following more intuitive equivalencies, which capture the notion of a simultaneous cutoff changes along relevant directions.
@@
Let $D$ be a cutoff rule over $\mathbb{R}^{K}$. A unit $i$ is
@@
The terminology established in Definition 3 indeed introduces well-defined categories:
@@
(b) Unit categorizations regarding (T,D)(T,\,D). While $\operatorname{supp}(T)$ captures change directions of $T$, the unit behavior is determined by $N^{T}$. The space is divided into (horizontally depicted) areas of alwaystaker, indecisives, complier, indecisives and nevertakers of $T$ with respect to $D$.
Figure 2: General case with $D$ not being a cutoff-rule. The decision boundaries of $T$ and $D$ are indicated with orange and blue dashed lines.
From now on, whenever we assume that $D$ is a cutoff rule, we restrict ourselves to the case in which $D$ has at least as much information for decision-making as $T$. This means that $D$ might depend on $I_{k,i}(0)$ for $k\in S(T)$. In addition, we suppose that $D$ does not depend on $I_{k,i}(c)$ with $c\neq 0$ for $k\in\operatorname{supp}(T)$, effectively restricting to the case where both decision rules have a zero cutoff. This restricts the general case. Suppose that the decision maker $D$ is an opportunist and ignores all the other scores as long as a certain incentive $X_{k}$ exceeds an even higher cutoff $0<c_{k}<X_{k}$. Only then $D$ would comply with $T$.
Even for cutoff rules $D$ with these restrictions, indecisive units are possible. For example let $D\coloneqq(\overline{I}_{1}\land I_{2})$ and $T\coloneqq I_{1}\land I_{2}$. Then $\operatorname{supp}(T)=\mathbb{R}^{2}=\operatorname{supp}(D)$ and for T(0| 0)=D(0| 0)T(0\,|\,0)=D(0\,|\,0) but $0=T(0\,|(-1,1))\neq D(0\,|\,(-1,1))=1$. Thus, $D$ is not constant nor equal to $T$ or $\overline{T}$ on the support of $T$. Given $T$ with $\dim(\operatorname{supp}(T))>1$, one can always construct a cutoff rule $D$ that exhibits indecisive items. At least one has the following:
@@
Moreover, our definitions of complier, nevertaker and alwaystaker imply the corresponding definitions in Imbens and Lemieux (2008). For this, let $c^{+},c^{-}\in\operatorname{supp}(T)$ be two directions that induce change in $T$, that is:
@@
Using the above complier definition:
@@
where $D_{i}(c)\coloneqq D(X_{i}-c)$ for $c\in\operatorname{supp}(T)$. The corresponding statements for nevertaker and alwaystaker follow analogously. Indeed, by requiring consistency in the limit for any direction, one would arrive at a local definition of the unit categories that is sufficient for the multi-score RDD setting. Before continuing with the identification part, we are going to expand on the AND-rule example.
@@
, dropping one of the mixed terms. Additionally, we have the following unit categorizations:
@@
    Note that $i$ being in the set on the right is equivalent to $\land_{j=k+1}^{K}I_{j,i}(c)$ being zero for all $c\in\operatorname{supp}(T)$. This is equivalent to $D_{i}(c)=0$ for all $c\in\operatorname{supp}(T)$, since given an $i$ one can always find a cutoff $c\in\operatorname{supp}(T)$ such that $T_{i}(c)=1$, which implies that one of the indicators $I_{j,i}$ for $j>k$ has to be zero.
@@
    Note that $T_{i}(c)=0$ implies $D_{i}(c)=0$, and that $c\in\operatorname{supp}(T)$ can always be chosen such that the former is satisfied.
Further examples can be found in the Appendix A.2 showing instances for the remaining unit categories (excluding the indecisive case). We conclude that the introduced definitions are indeed reasonable.
@@ 3.5 Effect Identification
Inspired by the work of Hahn et al. (2001); Imbens and Angrist (1994) we are going to use the introduced unit categories to prove an identification theorem for the effect on complier at the cutoff. For this section, we require that the outcome $Y$ does not directly depend on the treatment assignment $T$. Denote the set of all unit categories by
@@
This requirement relates to the independence assumptions used in Imbens and Angrist (1994). Using Assumption 2 and further assuming $\Pr\left(i\in\operatorname{Cat}\,\middle|\,X^{T}_{i}=0\right)>0$ one has
@@
We further rely on the following local continuity assumption, which is a variation of the standard continuity assumption (Assumption 1) adapted to the unit categories:
@@
This assumption can be weakened by assuming only directional continuity, which would render the effect dependent on the chosen directions. Note that
@@
holds for the non-change unit categories. Together with Assumption 3 one derives:
@@
Two more assumptions are required to make further use of the continuity. First, we deny the existence of indecisive items, since this category does not allow structured conclusions about $D$ based on knowledge of $T$. That is, this category does not separate the potential outcomes $Y_{i}(0)$ and $Y_{i}(1)$.
@@
Second, we assume that the directions $x^{+},\,x^{-}\in\operatorname{supp}(T)$ along which we aim to estimate the complier effect induce a change in $T$.
@@
This assumption is implicit in one-dimensional RDD designs and imposes no real restriction in practice, as $T$ is generally assumed to be known. With this, we know how $D_{i}$ behaves for complier and defier when approaching from $x^{+}$ and $x^{-}$ direction. That is:
@@
Thus, we can apply Assumption 3 to these two remaining categories on the right side of Equation 4 as well:
@@
with $C$ being the correction term for defier:
@@
This identification result allows for two immediate conclusions. First, one is free to choose among the directions $x^{+}$ and $x^{-}$ satisfying Assumption 5. Second, the proof suggests that dropping subsets $\Omega\subset\operatorname{Nt}\cup\operatorname{At}$ does not affect identification, as long as doing so does not violate Assumptions 3 and 2.
Following this thought we now investigate how dropping units in $\operatorname{At}\cup\operatorname{Nt}$ affects the above identification result. For ease of presentation, we assume that there do not exist any defier at the cutoff. As a consequence, the term $C$ in Equation 5 is equal to zero. Now let $\Omega\subset\operatorname{At}\cup\operatorname{Nt}$. Then $\Pr\left(i\in\operatorname{ComP},\,i\in\Omega\,\middle|\,X_{i}^{T}=0\right)=0$ and thus one has
@@
for the denominator in Equation 5. For the nominator,
@@
holds. Requiring Assumption 3 and Assumption 2 to hold when conditioning on $\Omega\cap\operatorname{Nt}$ and $\Omega\cap\operatorname{At}$ (instead of $\operatorname{Nt}$ and $\operatorname{At}$), we get:
@@
holds, we require an assumption similar to Assumption 2 for $\Pr\left(i\in\Omega\,\middle|\,X_{i}^{T}=\lambda x^{\pm}\right)$, in order to get
@@
Let Assumptions 2, 3, 4 and 5 hold, and $\Pr(i\in\operatorname{DeF}\,|\,X_{i}^{T}=0)=0$. Further, let $\Omega\subset\operatorname{Nt}\cup\operatorname{At}$ such that
@@
We call estimates of the complier effect of $T$ when removing $\Omega$ the sub-set complier effect of $T$ excluding $\Omega$. Theorem 2 provides sufficient conditions when both effects (the complier effect as identified in Theorem 1 and the sub-set complier effect) are equal.

### v2

@@ 3.3 Setup
We use the notation $\mathbb{R}^{K}_{>0}\coloneqq\left\{x\in\mathbb{R}^{K}\,\middle|\,\forall\,0\leq j\leq K\,:\,x_{j}>0\right\}$ for vectors with positive components and adapt it in the canonical way for the cross product of other ordered sets and relations (e.g., $\mathbb{R}^{K}_{<y}$ or $\mathbb{N}_{\leq k}$). Let $e_{1},\ldots,e_{K}\in\mathbb{R}^{K}$ denote the standard basis vectors, and let $\left\langle\cdot,\,\cdot\right\rangle$ the standard inner product in $\mathbb{R}^{K}$. For linear spaces $U$ and $V$ over $\mathbb{R}$ with $U\cap V=\{0\}$, we write $U\oplus V$ for their direct sum.
For each individual $i$, let $X_{i}=(X_{1,i},\,\ldots,\,X_{K,i})\in\mathbb{R}^{K}$ denote the vector of score variables. Given $c\in\mathbb{R}^{K}$, let $I_{i}(c)\coloneqq(I_{1,i}(c_{1}),\,\ldots,\,I_{K,i}(c_{K}))$ define the indicator vector, where $I_{k,i}(c_{k})\coloneqq\operatorname{\mathbbm{1}}[X_{k,i}>c_{k}]$. Let the observed outcome for individual $i$ be $Y_{i}$. We treat the entries of $I_{i}(c)$ as Boolean variables and allow any composition of AND, OR, and negation ($\land,\,\lor$$\overline{\square}$) over this set of atoms to form general Boolean functions $g(I_{i}(c))$.
@@
A mapping $T:\mathbb{R}^{K}\to\{0,1\}$ is called a decision rule. We say that a decision rule $T$ is a cutoff rule on $\mathbb{R}^{K}$ if there exists a Boolean mapping $g$ such that $T_{i}=T(X_{i})=g(I_{i})$. That is, $T$ determines the individual decisions $T_{i}$ by applying cutoffs to the components of $X_{i}$.
@@
holds for $\epsilon\in\mathbb{R}^{K},\lambda\in\mathbb{R}_{>0}$. Thus, without loss of generality, we assume the cutoff of interest is $c=0$. We suppose that the outcome $Y_{i}$ depends in the following way on a cutoff rule $T$ and a general decision rule $D$:
@@
where $T$ is the treatment assignment and $D$ is the implemented treatment. Unless otherwise stated, we impose no further assumptions on $D$, except that it is a decision rule.
@@ 3.4 Unit Categorization
Given this setup, certain groups of individuals are of particular interest: nevertakers, alwaystakers, compliers and defiers with respect to the pair (T,D)(T,\,D) or to (G,D)(G,\,D), where $G$ is a subrule of $T$. A meaningful categorization of unit $i$’s requires counterfactual reasoning: we must consider how hypothetical changes in the cutoff (or, equivalently, in the observed score values) would have changed the responses $T_{i}$ and $D_{i}$.
@@
Figure 1: Cutoff rules $D=I_{1}\land I_{2}$ and $T=I_{1}$. The decision boundaries coincide with the coordinate axes. When $X_{2}>0$, $T$ complies with $D$; otherwise, $D=0$ regardless of the value of $X_{1}$, which governs the behavior of $T$.
Figure 1 illustrates this idea for $D\coloneqq I_{1}\land I_{2}$ and $T\coloneqq I_{1}$, and shows that only changes $c\in\mathbb{R}^{2}$ that affect $T$ are relevant for categorizing the behavior of a unit $i$ with respect to (T,D)(T,\,D). Additional variables that affect only $D$ may not be controllable or even observable. The following definition formalizes the intuition behind relevant directions of change in a cutoff rule.
@@
Let $T$ be a cutoff rule on $\mathbb{R}^{K}$. Define
@@
as the set of support directions of $T$. The support of $T$ is then defined as the linear span of these directions:
@@
This technical definition greatly simplifies the treatment of equivalent cutoff rules for example when $T\coloneqq I_{1}\land(I_{2}\lor\overline{I_{2}})$ and $G\coloneqq I_{1}$. It also makes it possible to compare $T$ and $D$ on their common domain. The notion of the support of a cutoff rule is motivated by the existence of changes along coordinate directions. It follows almost immediately that if the support is empty, then no changes are possible along linear combinations of the coordinate directions. This property justifies the definition.
@@
From this point on, we assume that $T$ is non-degenerate, i.e., $\operatorname{supp}(T)\neq\{0\}$. In the introductory AND-rule example, with $D\coloneqq I_{1}(c)\land I_{2}(c)$ and $T\coloneqq I_{1}(c)$, we obtain $\operatorname{supp}(T)=\mathbb{R}\times\{0\}$ and $\operatorname{supp}(D)=\mathbb{R}^{2}$. If a unit $i$ complies with the assigned treatment $T$ relative to the actual treatment $D$, then $T_{i}$ should coincide with $D_{i}$ even under hypothetical changes of the cutoff $c_{1}$, as illustrated in Figure 1. A more refined, local definition could weaken this requirement to “any reasonable changes” (e.g. perturbations in a neighborhood of the cutoff), but for clarity we adopt the global one. In other words, both rules should agree on the support of $T$, which in this case occurs if $X_{2,i}>c_{2}$. Thus, it is $X_{2,i}$ and $c_{2}$ that govern the behavior of unit $i$. To make this distinction explicit, it is useful to introduce notation for entries $X\in\mathbb{R}^{K}$ that do not affect $T$:
@@
For each $X\in\mathbb{R}^{K}$, there exists a unique decomposition $X=X^{T}+X^{\perp T}$ with $X^{T}\in\operatorname{supp}(T)$ and $X^{\perp T}\in N^{T}$. The decomposition given by the orthogonal projection $P_{T}(X)\coloneqq\sum_{k\in S(T)}\left\langle X,\,e_{k}\right\rangle e_{k}$ onto $\operatorname{supp}(T)$ satisfies the above properties.
Thus, according to Proposition 1, the score space decomposes as
$\mathbb{R}^{|S(T)|}\times\mathbb{R}^{K-|S(T)|}\simeq\operatorname{supp}(T)\,\oplus N^{T}=\mathbb{R}^{K}.$
The first component captures the behavior of $T$, whereas the second component consists of free variables that do not affect $T$, but may influence the unit category. This observation motivates the following general definition of the unit categories.
@@
Let $i$ be a unit and let $X_{i}=X_{i}^{T}+X_{i}^{\perp T}$. Then $i$ is said to be:
@@
for all $c\in\operatorname{supp}(T)$. Let the sets of alwaystakers, nevertakers, compliers and defiers be denoted by $\operatorname{At}(T,D)$, $\operatorname{Nt}(T,D)$, $\operatorname{ComP}(T,D)$ and $\operatorname{DeF}(T,D)$.
In the special case where $D$ is a cutoff rule, these definitions admit the following intuitive equivalencies, which capture the notion of simultaneous cutoff changes along the relevant directions.
@@
Let $D$ be a cutoff rule on $\mathbb{R}^{K}$. The unit $i$ is:
@@
Definition 3 indeed introduces well-defined categories, which can be summarized as follows:
@@
(b) Unit categorizations with respect to (T,D)(T,\,D). While $\operatorname{supp}(T)$ captures the directions in which $T$ can change, unit behavior is determined by $N^{T}$. The space is partitioned (from top to bottom) into areas of alwaystakers, indecisives, compliers, indecisives and nevertakers of $T$ with respect to $D$.
Figure 2: General case with $D$ not being a cutoff rule. The decision boundaries of $T$ and $D$ are indicated with orange and blue dashed lines.
From this point forward, whenever we assume that $D$ is a cutoff rule, we restrict attention to the cases in which $D$ contains at least as much information for decision-making as $T$. This means that $D$ might depend on $I_{k,i}(0)$ for $k\in S(T)$. In addition, we suppose that $D$ does not depend on $I_{k,i}(c)$ with $c\neq 0$ for $k\in\operatorname{supp}(T)$, effectively restricting attention to rules with zero cutoff. This restriction narrows the general case. For example, suppose the decision-maker $D$ behaves opportunistically, ignoring all but one score $X_{k}$, only complying with $T$ when $X_{k}$ exceeds some higher cutoff $0<c_{k}<X_{k}$.
Even for cutoff rules $D$ under these restrictions, indecisive units are possible. For example, let $D\coloneqq(\overline{I}_{1}\land I_{2})$ and $T\coloneqq I_{1}\land I_{2}$. Then $\operatorname{supp}(T)=\mathbb{R}^{2}=\operatorname{supp}(D)$ and for T(0| 0)=D(0| 0)T(0\,|\,0)=D(0\,|\,0) but $0=T(0\,|(-1,1))\neq D(0\,|\,(-1,1))=1$. Thus, $D$ is not constant and differs from $T$ and $\overline{T}$ on the support of $T$. Given $T$ with $\dim(\operatorname{supp}(T))>1$, one can always construct a cutoff rule $D$ that produces indecisive items. At least one can prove the following:
@@
Moreover, our definitions of alwaystakers, nevertakers and compliers imply the corresponding definitions in Imbens and Lemieux (2008). For this, let $c^{+},c^{-}\in\operatorname{supp}(T)$ be two directions that induce change in $T$, so that:
@@
Using the above complier definition
@@
where $D_{i}(c)\coloneqq D(X_{i}-c)$ for $c\in\operatorname{supp}(T)$. The corresponding statements for nevertakers and alwaystakers follow analogously. By requiring consistency in the limit for any direction, one arrives at a local definition of unit categories that is sufficient for the multi-score RDD setting. Before turning to the identification results, we expand on the AND-rule example.
@@
, dropping one of the mixed terms. Additionally, the unit categorizations are as follows:
@@
    A unit $i$ being in the set on the right side is equivalent to $\land_{j=k+1}^{K}I_{j,i}(c)$ being zero for all $c\in\operatorname{supp}(T)$. This is equivalent to $D_{i}(c)=0$ for all $c\in\operatorname{supp}(T)$, since for this unit one can always find a cutoff $c\in\operatorname{supp}(T)$ such that $T_{i}(c)=1$, which implies that one of the indicators $I_{j,i}$ for $j>k$ has to be zero.
@@
    Note that $T_{i}(c)=0$ implies $D_{i}(c)=0$ for $c\in\operatorname{supp}(T)$. Now suppose that $T_{i}(c)=1$ and $D_{i}(c)=0$ for some $c\in\operatorname{supp}(T)$. Thus, $I_{j,i}(c)=0$ for some $j>k$. Since $I_{j,i}$ independent of $c\in\operatorname{supp}(T)$, $D_{i}=0$ on $\operatorname{supp}(T)$.
Further examples can be found in Appendix A.2, which introduces additional instances of the remaining unit categories (excluding the indecisive case). We conclude that the definitions introduced above are reasonable.
@@ 3.5 Effect Identification
Inspired by the work of Hahn et al. (2001) and Imbens and Angrist (1994), we use the unit categories introduced above to prove an identification theorem for the local complier effect at the cutoff. Throughout this section, we require that the outcome $Y$ does not directly depend on the treatment assignment $T$. Let the set of all unit categories be
@@
This assumption relates to the independence assumptions in Imbens and Angrist (1994). Using Assumption 2 and further assuming $\Pr\left(i\in\operatorname{Cat}\,\middle|\,X^{T}_{i}=0\right)>0$ we obtain
@@
We also rely on the following local continuity assumption, which adapts the standard continuity condition (Assumption 1) to the unit categories:
@@
This assumption can be weakened by requiring only directional continuity, in which case the effect would depend on the chosen directions. Note that
@@
holds for the non-change unit categories. Together with Assumption 3 this yields:
@@
Two further assumptions are needed to make use of the continuity condition for the remaining categories. First, we rule out the existence of indecisive units, since this category does not allow structured conclusions about $D$ based on knowledge of $T$. In other words, this category does not allow separating the potential outcomes $Y_{i}(0)$ and $Y_{i}(1)$.
@@
Second, we assume that the directions $x^{+},\,x^{-}\in\operatorname{supp}(T)$ along which we estimate the complier effect induce a change in $T$.
@@
This assumption is implicit in one-dimensional RDD designs and imposes no substantive restriction in practice, as $T$ is typically known. With this in place, we know how $D_{i}$ behaves for compliers and defiers when approaching from the $x^{+}$ and $x^{-}$ directions. That is:
@@
Thus, we can apply Assumption 3 to these two remaining categories on the right-hand side of Equation 4 as well:
@@
with $C$ being the correction term for defiers:
@@
This identification result has two immediate implications. First, one is free to choose among the directions $x^{+}$ and $x^{-}$ satisfying Assumption 5. Second, the proof suggests that dropping subsets $\Omega\subset\operatorname{Nt}\cup\operatorname{At}$ does not affect identification, as long as doing so does not violate Assumptions 3 and 2.
Following this idea, we now investigate how excluding units in $\operatorname{At}\cup\operatorname{Nt}$ affects the above identification result. For ease of presentation, we assume there are no defiers at the cutoff. As a consequence, the correction term $C$ in Equation 5 equals zero. Now let $\Omega\subset\operatorname{At}\cup\operatorname{Nt}$. Then $\Pr\left(i\in\operatorname{ComP},\,i\in\Omega\,\middle|\,X_{i}^{T}=0\right)=0$ and thus one has
@@
for the denominator in Equation 5. For the numerator, we have:
@@
Requiring Assumptions 3 and 2 to hold when conditioning on $\Omega\cap\operatorname{Nt}$ and $\Omega\cap\operatorname{At}$ (instead of $\operatorname{Nt}$ and $\operatorname{At}$), we obtain:
@@
holds, we require an assumption similar to Assumption 2 for $\Pr\left(i\in\Omega\,\middle|\,X_{i}^{T}=\lambda x^{\pm}\right)$, in order to obtain
@@
Let Assumptions 2, 3, 4 and 5 hold, and suppose $\Pr(i\in\operatorname{DeF}\,|\,X_{i}^{T}=0)=0$. Further, let $\Omega\subset\operatorname{Nt}\cup\operatorname{At}$ such that:
@@
We refer to estimates of the complier effect of $T$ obtained when removing $\Omega$ as the subset complier effect of $T$ (excluding $\Omega$). Theorem 2 provides sufficient conditions when both effects (the complier effect as identified in Theorem 1 and the subset complier effect) are equal.